# Non-asymptotic Estimates for Markov Transition Matrices with Rigorous Error Bounds

### v2

@@ Abstract.
Markov models are fundamental in various fields for their efficiency in modeling time-dependent scenarios where the future state depends solely on the present state. The inverse problem of estimating the transition matrix of a given Markov chain is one of the most fundamental tasks in many scientific applications [18, 6, 12, 21, 17].
@@
Then $\mu(u)$ and $p(u,v)$ can be approximated by $\sum_{w\in\Omega}r_{n}(u,w)$ and $r_{n}(u,v)/\sum_{w\in\Omega}r_{n}(u,w)$, respectively. We will use $\bm{R}_{n}=[r_{n}(u,v)]_{u,v\in\Omega}$ to denote the matrix form of MLE. It is shown in [1, 3] that the error of MLE is asymptotically $\mathcal{O}(1/\sqrt{n})$ by the central limiting theorem for Markov chains. In contrast, our goal is to establish non-asymptotic results that are valid for any $n\in\mathbb{N}^{+}$.
Our first contribution is to establish non-asymptotic error bounds of MLE in terms of the iterated Poincaré gap (IP gap) of $\bm{P}$, which was first introduced in [5] as a generalization of the spectral gap for non-reversible Markov chains. We call it the iterated Poincaré gap gap, following the notations in [10], to emphasize its definition by an iterated Poincaré gap inequality (see Subsection 2.1 below). To achieve this, we interpret MLE as the Markov empirical mean of the indicator function $\mathbb{1}_{(u,v)}(\cdot,\cdot)$ corresponding to some transition matrix $\bm{P}_{2}$ defined on the length-2 path space $\Omega^{2}$. We then establish the relation between the spectral gaps of $\bm{P}_{2}$ and $\bm{P}$, and apply concentration inequalities for Markov chains to derive the following entry-wise tail bound for the error of MLE.
@@
There have been many related works that also focus on learning and testing Markov transition matrices [11, 23, 4, 25]. Most relevantly, [23] controls the probabilistic sample complexity of MLE at the matrix level (in terms of the total variation norm). We will discuss their results in detail in Section 2. We want to emphasize that, compared with prior work, the novelty of our result lies in that it only requires a nonzero IP gap of $\bm{P}$ (which proves to be more robust than other generalized spectral gaps for non-reversible Markov chains; see [5, 10]), owing to a recently established Bernstein-type inequality [10] by the same authors. We will show that $\eta_{p}(\bm{P})>0$ holds for any irreducible transition matrix $\bm{P}$. Hence, Theorem 1.1 applies to all Markov chains with irreducible transition matrices, highlighting the robustness of our result.
In numerous applications, the Markov chain considered is reversible [9], i.e. the transition matrix satisfies the detailed balance condition. In this case, it is important that the estimator preserves the reversibility of $\bm{P}$  [2, 21, 17]. However, the theories proposed in the literature do not provide any non-asymptotic convergence rate guarantee for those estimators.
@@
Let $\tilde{\bm{P}}_{2}$ denote the corresponding transition matrix on $\tilde{\Omega}^{2}$. A non-asymptotic error bound of SCE in the matrix form can be obtained by exploiting the spectral properties of $\tilde{\bm{P}}_{2}$ and applying concentration inequalities for Markov dependent random matrices recently established in [19].
@@ 2.1.1. Notation
Throughout this paper, matrices are written in uppercase bold letters. In particular, $\bm{E}_{uv}$ denotes the matrix with $1$ at position $(u,v)$ and 00 elsewhere, $\bm{\mathbf{I}}_{d}$ denotes the $d$-dimensional identity matrix, and $\bm{1}_{d}$ denotes the d-dimensional all-one vector. In cases where it does not cause ambiguity, we will omit the subscript $d$ to use $\bm{\mathbf{I}}$ and $\bm{1}$ instead of $\bm{\mathbf{I}}_{d}$ and $\bm{1}_{d}$.
@@ 2.1.2. Foundation of Markov chains
This definition coincides with that of the non-reversible spectral gap first formally introduced by Chatterjee in [5], where he used this quantity to establish ergodicity for non-reversible Markov chains, generalizing similar results for reversible Markov chains based on the usual spectral gap. The reader may thus also refer to the IP gap as the Chatterjee gap or the non-reversible spectral gap and can find a more comprehensive introduction to the topic in [5]. This quantity was independently developed by [10] to establish Bernstein-type inequalities for Markov chains. In this paper, we choose to adopt the notation in [10] and call this quantity the iterated Poincaré gap, since it is related to the iterated Poincaré inequality:
@@ 2.3. Generalized spectral gaps
In the non-reversible case, various notions of spectral gaps have been proposed as substitutes for the usual spectral gap before [5]. A commonly used quantity is the absolute $L_{2}$-spectral gap (or simply the absolute spectral gap), which is defined as $\eta_{a}(\bm{P})=1-\lambda_{a}(\bm{P})$, where
@@
Here $\bm{\Pi}:=\bm{1}\mu^{\mathrm{T}}$ is the projection onto $\bm{1}$ under the $L_{2,\mu}$ inner product. One can see that $\eta_{a}(\bm{P})$ is the gap between $1$ and the second largest singular value of $\bm{P}$ under the $L_{2,\mu}$ inner product. Alternatively, $\eta_{a}(\bm{P})$ can be understood as the usual spectral gap of the self-adjoint operator $\left(\bm{P}^{*}\bm{P}\right)^{1/2}$, where $\bm{P}^{*}$ is the adjoint of $\bm{P}$ under the $L_{2,\mu}$ inner product. This quantity was widely used in the literature to bound the mixing time [7] and establish concentration inequalities [14, 15] for non-reversible Markov chains.
A generalization of the absolute spectral gap, called the pseudo spectral gap, was first introduced in [20] to establish concentration inequalities for non-reversible Markov chains. It is defined as
@@
Advanced theoretical studies, estimations and applications of the pseudo spectral gap can be found in [22, 24].
@@
It is shown in [7] that $\eta_{s}(\bm{P})$ can be used to get upper bounds on the mixing time of $\bm{P}$ when $\bm{P}$ is strongly aperiodic. Moreover, the additive reversiblization $\bm{A}$ is usually more user-friendly than the the multiplicative reversiblization $\bm{P}^{*}\bm{P}$. However, except for [8, 13], which appear to contain a technical gap in the proof of a key lemma (see [10] for details), there are no concentration results that depend on $\eta_{s}(\bm{P})$.
@@ 2.4. Related works
There have been many related works that also focus on learning and testing Markov transition matrices [11, 23, 4, 25]. Most relevantly, [23] controls the probabilistic sample complexity of MLE in terms of the mixing time $t_{\text{mix}}$. More specifically, we say the transition matrix $\bm{P}$ is geometrically ergodic, if there exists $\rho\in(0,1)$ such that for any $u\in\Omega$,
@@
It is shown in [23, Theorem 3.3] that if $\bm{P}$ is geometrically ergodic, then for any $\varepsilon,\delta>0$ and any
@@
It is worth-noting that [23, Theorem 3.3] only works when the transition matrix $\bm{P}$ is geometrically ergodic. However, by the central limiting theorem for Markov chains, one does not need the geometrically ergodicity of $\bm{P}$ to guarantee the convergence of the Markov empirical mean to its space average. In comparison, our Theorem 2.1(2) establishes probabilistic error bound for MLE at the entry-wise level, which can be applied to any irreducible Markov chains (not necessarily geometrically ergodic). Note that $r_{n}(u,v)$ can be interpreted as the Markov empirical mean of the indicator function $\mathbb{1}_{(u,v)}(\cdot,\cdot)$ corresponding to $\bm{P}_{2}$. Therefore, a recently established Bernstein-type inequality for Markov sums [10], under milder assumptions than [20, Corollary 2.10], can be applied to $\bm{P}_{2}$ to derive that
@@
One may also consider applying typical concentration inequalities for Markov sums [14, 15, 16, 20, 13, 8] to establish tail bounds similar to Theorem 2.1(2). This will require analyzing other spectral gaps of $\bm{P}_{2}$ defined in Subsection 2.3, which wil also be presented in Theorem 3.3 as a result of independent interest. Specifically, we show that $\eta_{a}(\bm{P}_{2})\equiv 0$ for any transition matrix $\bm{P}$. Hence, most typical concentration results are inapplicable to to $\bm{P}_{2}$, except for [20, Theorem 3.4]. On the other hand, we prove that $\eta_{ps}(\bm{P}_{2})\geq\eta_{ps}(\bm{P})/2$, which allows us to apply [20, Theorem 3.4] to deduce that
@@
if $n\gtrsim 1/(\eta_{ps}(\bm{P}))$. However, as shown in Subsection 2.3, if $\bm{P}$ is not geometrically ergodic, then it is possible that $\eta_{ps}(\bm{P})=0$. Moreover, since $\eta_{p}(\bm{P})\geq\eta_{ps}(\bm{P})/2$, the bound in (2.17) does not provide any improvement over the spectral gap term of Theorem 2.1(2). In fact, since all other spectral gaps introduced in Subsection 2.3 can be bounded in terms of the IP gap, the only potential benefit of employing other concentration results instead of [10] lies in achieving a possibly better constant $c$ in the resulting tail bound.
Since Theorem 2.1(2) is an entry-wise result, it is natural to ask if we can establish a probabilistic error bound for MLE at the matrix level that depends on the IP gap of $\bm{P}$. The main challenge in obtaining a result analogous to [23, Theorem 3.3] lies in studying the concentration of the error norm (2.16), since Paulin’s result [20, Corollary 2.10] requires a finite mixing time, which is not guaranteed for all irreducible transition matrices. It would be interesting to develop a Markov concentration inequality for general functions of $n$ variables instead of sums based on the IP gap. If such a result is established, one could then derive a matrix-level bound analogous to [23, Theorem 3.3]. We will leave it as a future work.
On the other hand, when $\bm{P}$ is reversible, the transition matrix $\tilde{\bm{P}}_{2}$ defined on the symmetric path space $\tilde{\Omega}^{2}$ always admits a finite mixing time, which can be shown to be bounded by the mixing time of $(\bm{P}+\bm{I})/2$ plus 1 similar to [23, Lemma 6.1]. Therefore, following the proof strategy similar to [23, Theorem 3.3], one can derive a matrix-level bound for SCE in terms of the mixing time of $(\bm{P}+\bm{I})/2$, which can further be bounded by $c\log(\mu_{*})/\eta(\bm{P})$. Here $\mu_{*}:=\min_{u\in\Omega}\mu(u)>0$ by the Perron–Frobenius theorem. We do not pursue this approach here, as the proof is essentially the same. Instead, we establish Theorem 2.2(2) by applying a recently established matrix concentration inequality [19] to $\tilde{\bm{P}}_{2}$, providing a matrix-level error bound for SCE via a more direct method. Note, however, that the result in [19] requires a nonzero absolute spectral gap of the transition matrix, which does not hold for $\bm{P_{2}}$. Therefore, extending this approach to MLE would require a matrix concentration result based on more robust spectral gaps (e.g., the IP gap), which is another interesting direction for future research.
@@ 3.1.1. Construction of $\bm{P}_{2}$
Recall that $\bm{P}=[p(u,v)]_{u,v\in\Omega}$ is a transition matrix on the state space $\Omega$ with an unique invariant distribution $\mu$ Then the length-2 path space $\Omega^{2}=\{(u_{1},u_{2}):u_{1},u_{2}\in\Omega\}$ is naturally equipped with a distribution $\mu_{2}$ that is given by
@@ 3.1.2. Construction of $\tilde{\bm{P}}_{2}$
    {choose $\mathbb{P}\{\text{choose }v\}=p\left(w_{2},v\right),$ (3.7)
@@ 3.1.3. Invariant distributions
$\tilde{\mu}_{2}(\tilde{\bm{u}})\tilde{p}_{2}(\tilde{\bm{u}},\tilde{\bm{v}})=\mu(u)p(u,w)p(w,v)=\mu(v)p(v,w)p(w,u)=\tilde{\mu}_{2}(\tilde{\bm{v}})\tilde{p}_{2}(\tilde{\bm{v}},\tilde{\bm{u}}),$ (3.11)
@@ 3 $\hat{\mu}=\bm{R}\bm{1}$;
Input: An oracle $\mathcal{P}$ of a reversible Markov chain, initial distribution $\nu$, total iteration number $N$;

### v3

# Non-asymptotic Estimates for Markov Transition Matrices via Spectral gap methods

@@ Abstract.
Markov models are fundamental in various fields for their efficiency in modeling time-dependent scenarios where the future state depends solely on the present state. The inverse problem of estimating the transition matrix of a given Markov chain is one of the most fundamental tasks in many scientific applications [mostel2020statistical, craig2002estimation, inamura2006estimating, trendelkamp2015estimation, macdonald2021fitting].in
@@
Then $\mu(u)$ and $p(u,v)$ can be approximated by $\sum_{w\in\Omega}r_{n}(u,w)$ and $r_{n}(u,v)/\sum_{w\in\Omega}r_{n}(u,w)$, respectively. We will use $\bm{R}_{n}=[r_{n}(u,v)]_{u,v\in\Omega}$ to denote the matrix form of MLE. It is shown in [anderson1957statistical, billingsley1961statistical] that the error of MLE is asymptotically $\mathcal{O}(1/\sqrt{n})$ by the central limiting theorem for Markov chains. In contrast, our goal is to establish non-asymptotic results that are valid for any $n\in\mathbb{N}^{+}$.
Our first contribution is to establish non-asymptotic error bounds of MLE in terms of the iterated Poincaré gap (IP gap) of $\bm{P}$, which was first introduced in [chatterjee2025spectral] as a generalization of the spectral gap for non-reversible Markov chains. We call it the iterated Poincare gap, following the notations in [huang2024bernsteintypeinequalitiesmarkovchains], to emphasize its definition by an iterated Poincare inequality (see Subsection 2.1 below). To achieve this, we interpret MLE as the Markov empirical mean of the indicator function $\mathbb{1}_{(u,v)}(\cdot,\cdot)$ corresponding to some transition matrix $\bm{P}_{2}$ defined on the length-2 path space $\Omega^{2}$. We then establish the relation between the spectral gaps of $\bm{P}_{2}$ and $\bm{P}$, and apply concentration inequalities for Markov chains to derive the following entry-wise tail bound for the error of MLE.
@@
There have been many related works that also focus on learning and testing Markov transition matrices [hao2018learning, wolfer2021statistical, chan2021learning, wolfer2024empirical]. Most relevantly, [wolfer2021statistical] controls the probabilistic sample complexity of MLE at the matrix level (in terms of the total variation norm). We will discuss their results in detail in Section 2. We want to emphasize that, compared with prior work, the novelty of our result lies in that it only requires a nonzero IP gap of $\bm{P}$ (which proves to be more robust than other generalized spectral gaps for non-reversible Markov chains; see [chatterjee2025spectral, huang2024bernsteintypeinequalitiesmarkovchains]), owing to a recently established Bernstein-type inequality [huang2024bernsteintypeinequalitiesmarkovchains] by the same authors. We will show that $\eta_{p}(\bm{P})>0$ holds for any irreducible transition matrix $\bm{P}$. Hence, Theorem 1.1 applies to all Markov chains with irreducible transition matrices, highlighting the robustness of our result.
In numerous applications, the Markov chain considered is reversible [grone2008interlacing], i.e. the transition matrix satisfies the detailed balance condition. In this case, it is important that the estimator preserves the reversibility of $\bm{P}$  [annis2010estimation, trendelkamp2015estimation, macdonald2021fitting]. However, the theories proposed in the literature do not provide any non-asymptotic convergence rate guarantee for those estimators.
@@
Let $\tilde{\bm{P}}_{2}$ denote the corresponding transition matrix on $\tilde{\Omega}^{2}$. A non-asymptotic error bound of SCE in the matrix form can be obtained by exploiting the spectral properties of $\tilde{\bm{P}}_{2}$ and applying concentration inequalities for Markov dependent random matrices recently established in [neeman2024concentration].
@@ 2.1.1. Notation
Throughout this paper, matrices are written in uppercase bold letters. In particular, $\bm{E}_{uv}$ denotes the matrix with $1$ at position $(u,v)$ and 0 elsewhere, $\bm{\mathbf{I}}_{d}$ denotes the $d$-dimensional identity matrix, and $\bm{1}_{d}$ denotes the d-dimensional all-one vector. In cases where it does not cause ambiguity, we will omit the subscript $d$ to use $\bm{\mathbf{I}}$ and $\bm{1}$ instead of $\bm{\mathbf{I}}_{d}$ and $\bm{1}_{d}$.
@@ 2.1.2. Foundation of Markov chains
This definition coincides with that of the non-reversible spectral gap first formally introduced by Chatterjee in [chatterjee2025spectral], where he used this quantity to establish ergodicity for non-reversible Markov chains, generalizing similar results for reversible Markov chains based on the usual spectral gap. The reader may thus also refer to the IP gap as the Chatterjee gap or the non-reversible spectral gap and can find a more comprehensive introduction to the topic in [chatterjee2025spectral]. This quantity was independently developed by [huang2024bernsteintypeinequalitiesmarkovchains] to establish Bernstein-type inequalities for Markov chains. In this paper, we choose to adopt the notation in [huang2024bernsteintypeinequalitiesmarkovchains] and call this quantity the iterated Poincaré gap, since it is related to the iterated Poincaré inequality:
@@ 2.3. Generalized spectral gaps
In the non-reversible case, various notions of spectral gaps have been proposed as substitutes for the usual spectral gap before [chatterjee2025spectral]. A commonly used quantity is the absolute $L_{2}$-spectral gap (or simply the absolute spectral gap), which is defined as $\eta_{a}(\bm{P})=1-\lambda_{a}(\bm{P})$, where
@@
Here $\bm{\Pi}:=\bm{1}\mu^{\mathrm{T}}$ is the projection onto $\bm{1}$ under the $L_{2,\mu}$ inner product. One can see that $\eta_{a}(\bm{P})$ is the gap between $1$ and the second largest singular value of $\bm{P}$ under the $L_{2,\mu}$ inner product. Alternatively, $\eta_{a}(\bm{P})$ can be understood as the usual spectral gap of the self-adjoint operator $\left(\bm{P}^{*}\bm{P}\right)^{1/2}$, where $\bm{P}^{*}$ is the aadjoint of $\bm{P}$ under the $L_{2,\mu}$ inner product. This quantity was widely used in the literature to bound the mixing time [fill1991eigenvalue] and establish concentration inequalities [lezaud1998chernoff, leon2004optimal] for non-reversible Markov chains.
A generalization of the absolute spectral gap, called the pseudo spectral gap, was first introduced in [paulin2015concentration] to establish concentration inequalities for non-reversible Markov chains. It is defined as
@@
Advanced theoretical studies, estimations and applications of the pseudo spectral gap can be found in [wolfer2019estimating, wolfer2024improved].
@@
It is shown in [fill1991eigenvalue] that $\eta_{s}(\bm{P})$ can be used to get upper bounds on the mixing time of $\bm{P}$ when $\bm{P}$ is strongly aperiodic. Moreover, the additive reversiblization $\bm{A}$ is usually more user-friendly than the the multiplicative reversiblization $\bm{P}^{*}\bm{P}$. However, except for [fan2021hoeffding, jiang2018bernstein], which appear to contain a technical gap in the proof of a key lemma (see [huang2024bernsteintypeinequalitiesmarkovchains] for details), there are no concentration results that depend on $\eta_{s}(\bm{P})$.
@@ 2.4. Related works
There have been many related works that also focus on learning and testing Markov transition matrices [hao2018learning, wolfer2021statistical, chan2021learning, wolfer2024empirical]. Most relevantly, [wolfer2021statistical] controls the probabilistic sample complexity of MLE in terms of the mixing time $t_{\text{mix}}$. More specifically, we say the transition matrix $\bm{P}$ is geometrically ergodic, if there exists $\rho\in(0,1)$ such that for any $u\in\Omega$,
@@
It is shown in [wolfer2021statistical, Theorem 3.3] that if $\bm{P}$ is geometrically ergodic, then for any $\varepsilon,\delta>0$ and any
@@
It is worth-noting that [wolfer2021statistical, Theorem 3.3] only works when the transition matrix $\bm{P}$ is geometrically ergodic. However, by the central limiting theorem for Markov chains, one does not need the geometrically ergodicity of $\bm{P}$ to guarantee the convergence of the Markov empirical mean to its space average. In comparison, our Theorem 2.1(2) establishes probabilistic error bound for MLE at the entry-wise level, which can be applied to any irreducible Markov chains (not necessarily geometrically ergodic). Note that $r_{n}(u,v)$ can be interpreted as the Markov empirical mean of the indicator function $\mathbb{1}_{(u,v)}(\cdot,\cdot)$ corresponding to $\bm{P}_{2}$. Therefore, a recently established Bernstein-type inequality for Markov sums [huang2024bernsteintypeinequalitiesmarkovchains], under milder assumptions than [paulin2015concentration, Corollary 2.10], can be applied to $\bm{P}_{2}$ to derive that
@@
One may also consider applying typical concentration inequalities for Markov sums [lezaud1998chernoff, leon2004optimal, miasojedow2014hoeffding, paulin2015concentration, jiang2018bernstein, fan2021hoeffding] to establish tail bounds similar to Theorem 2.1(2). This will require analyzing other spectral gaps of $\bm{P}_{2}$ define in Subsection 2.3, which wil also be presented in Theorem 3.3 as a result of independent interest. Specifically, we show that $\eta_{a}(\bm{P}_{2})\equiv 0$ for any transition matrix $\bm{P}$. Hence, most typical concentration results are inapplicable to to $\bm{P}_{2}$, except for [paulin2015concentration, Theorem 3.4]. On the other hand, we prove that $\eta_{ps}(\bm{P}_{2})\geq\eta_{ps}(\bm{P})/2$, which allows us to apply [paulin2015concentration, Theorem 3.4] to deduce that
@@
if $n\gtrsim 1/(\eta_{ps}(\bm{P}))$. However, as shown in Subsection 2.3, if $\bm{P}$ is not geometrically ergodic, then it is possible that $\eta_{ps}(\bm{P})=0$. Moreover, since $\eta_{p}(\bm{P})\geq\eta_{ps}(\bm{P})/2$, the bound in (2.17) does not provide any improvement over the spectral gap term of Theorem 2.1(2). In fact, since all other spectral gaps introduced in Subsection 2.3 can be bounded in terms of the IP gap, the only potential benefit of employing other concentration results instead of [huang2024bernsteintypeinequalitiesmarkovchains] lies in achieving a possibly better constant $c$ in the resulting tail bound.
Since Theorem 2.1(2) is an entry-wise result, it is natural to ask if we can establish a probabilistic error bound for MLE at the matrix level that depends on the IP gap of $\bm{P}$. The main challenge in obtaining a result analogous to [wolfer2021statistical, Theorem 3.3] lies in studying the concentration of the error norm (2.16), since Paulin’s result [paulin2015concentration, Corollary 2.10] requires a finite mixing time, which is not guaranteed for all irreducible transition matrices. It would be interesting to develop a Markov concentration inequality for general functions of $n$ variables instead of sums based on the IP gap. If such a result is established, one could then derive a matrix-level bound analogous to [wolfer2021statistical, Theorem 3.3]. We will leave it as a future work.
On the other hand, when $\bm{P}$ is reversible, the transition matrix $\tilde{\bm{P}}_{2}$ defined on the symmetric path space $\tilde{\Omega}^{2}$ always admits a finite mixing time, which can be shown to be bounded by the mixing time of $(\bm{P}+\bm{I})/2$ plus 1 similar to [wolfer2021statistical, Lemma 6.1]. Therefore, following the proof strategy similar to [wolfer2021statistical, Theorem 3.3], one can derive a matrix-level bound for SCE in terms of the mixing time of $(\bm{P}+\bm{I})/2$, which can further be bounded by $c\log(\mu_{*})/\eta(\bm{P})$. Here $\mu_{*}:=\min_{u\in\Omega}\mu(u)>0$ by the Perron–Frobenius theorem. We do not pursue this approach here, as the proof is essentially the same. Instead, we establish Theorem 2.2(2) by applying a recently established matrix concentration inequality [neeman2024concentration] to $\tilde{\bm{P}}_{2}$, providing a matrix-level error bound for SCE via a more direct method. Note, however, that the result in [neeman2024concentration] requires a nonzero absolute spectral gap of the transition matrix, which does not hold for $\bm{P_{2}}$. Therefore, extending this approach to MLE would require a matrix concentration result based on more robust spectral gaps (e.g., the IP gap), which is another interesting direction for future research.
@@ 3.1.1. Construction of $\bm{P}_{2}$
Recall that $\bm{P}=[p(u,v)]_{u,v\in\Omega}$ is a transition matrix on the state space $\Omega$ with an unique invariant distribution $\mu$ . Then the length-2 path space $\Omega^{2}=\{(u_{1},u_{2}):u_{1},u_{2}\in\Omega\}$ is naturally equipped with a distribution $\mu_{2}$ that is given by
@@ 3.1.2. Construction of $\tilde{\bm{P}}_{2}$
    $\mathbb{P}\{\text{choose }v\}=p\left(w_{2},v\right),$ (3.7)
@@ 3.1.3. Invariant distributions
[…]
@@ 3 $\hat{\mu}=\bm{R}\bm{1}$;
Input: An oracle $\mathcal{P}$ of a reversible Markov chain, initial distribution $\nu$, total iteration number NN;